OpenStudy (anonymous):

how do you solve (n+1!)/n!

OpenStudy (anonymous):

n+1

OpenStudy (anonymous):

how

OpenStudy (anonymous):

(n + 1)! = (n+1) * n!

OpenStudy (anonymous):

the n! will cancel out

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

k thx :) then how would i do (2n-1)!/(2n+1)!

OpenStudy (anonymous):

(2n+1)! = (2n +1) (2n) (2n - 1)! answer is 1/(2n + 1)(2n)

OpenStudy (anonymous):

how about (2n+2)!/(2n)!

OpenStudy (anonymous):

(2n + 2) (2n + 1) (2n)!/(2n)! = (2n + 2) (2n + 1)

OpenStudy (anonymous):

can you please explain to me like the rules of n! and how or what order I can expand it in order to simpilfy?

OpenStudy (anonymous):

do u expand on the numberitor or demoninator/

OpenStudy (anonymous):

moneybird? plz help